Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability1

Expert replies
by gunjan1208 » Wed Oct 05, 2011 12:39 am
Please help me explaning the answer (Source MGMAT strategy guides)

A florist has 2 azaleas, 3 buttercups, and 4 petunias. She puts two flowers together at
random in a bouquet. However, the customer calls and says that she does not want
two of the same flower. What is the probability that the florist does not have to
change the bouquet?
Join the discussion
Source: — Problem Solving |

by sl750 » Wed Oct 05, 2011 12:52 am
P(2 azaleas) = 2/9C2 ; P(2 Buttercup) = 3/9C2; P(2 Petunias) = 4C2/9C2

1 - P(Bouquet with identical flowers) = P(bouquet having 2 different flowers)

2/72+6/72+12/72 = 20/72

1-5/18 = 13/18
Last edited by sl750 on Wed Oct 05, 2011 10:25 am, edited 1 time in total.
Join the discussion

by svd.kumar » Wed Oct 05, 2011 10:11 am
Please post the correct answer, I got a different result. Please correct me if I am wrong

2 azaleas, 3 buttercups, and 4 petunias

Probability of selecting 2 azaleas = 2/9*1/8
Probability of selecting 3 buttercups = 3/9*2/8
Probability of selecting 4 petunias = 4/9*3/8

Which gives 2/72 + 6/72 + 12/72 = 5/18

But we want probability of not having exactly same = 1 - 5/18 = 13/18

Cheers,
SVD
Join the discussion

by sl750 » Wed Oct 05, 2011 10:23 am
I guess I screwed up. 2C2 = 1
Join the discussion

by Anurag@Gurome » Wed Oct 05, 2011 8:03 pm
gunjan1208 wrote:Please help me explaning the answer (Source MGMAT strategy guides)

A florist has 2 azaleas, 3 buttercups, and 4 petunias. She puts two flowers together at
random in a bouquet. However, the customer calls and says that she does not want
two of the same flower. What is the probability that the florist does not have to
change the bouquet?
Let azaleas = A, buttercups = B, and petunias = P

Required probability = 1 - [Probability of both being A + probability of both being B + probability of both being P]
Probability of choosing the same flower = 2/9 * 1/8 + 3/9 * 2/8 + 4/9 * 3/8 = 5/18

Therefore, required probability = 1 = 5/18 = [spoiler]13/18[/spoiler]
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by tpr-becky » Wed Oct 05, 2011 8:05 pm
Probability at its most basic is want over total - so eventually we want to solve for

(bouquet without two same)/ All bouquets

First solve for how many bouquets can you make - with two flowers you have 9 to choose from for the first slot and 8 to choose from in the second slot but order doesn't matter so you divide by 2. Thus 9*8/2 = 36 total bouquets.

Now we need to find the nubmer of bouquets that won't have the same - there ways to look at this
Bouquet has an azalea and another flower (2 x 7)/2 = 7 bouquets
or
Bouquet has a buttercup and another flower: (3 x 6)/2 = 9 bouquets
or
Bouquet has a petunia and another flower: (4 x 5)/2 = 10 bouquets.

add up all bouquets that don't have double flowers - 26 options.

Now put you want over your total - for 26/36 = 13/18 is the final answer
Becky
Master GMAT Instructor
The Princeton Review
Irvine, CA
Join the discussion

by gunjan1208 » Thu Oct 06, 2011 2:35 am
Probability of choosing the same flower = 2/9 * 1/8 + 3/9 * 2/8 + 4/9 * 3/8 = 5/18


Why do we multiply with these boldenend numbers?
Join the discussion

by catfreak » Thu Oct 06, 2011 7:57 am
gunjan1208 wrote:Probability of choosing the same flower = 2/9 * 1/8 + 3/9 * 2/8 + 4/9 * 3/8 = 5/18


Why do we multiply with these boldenend numbers?
We are choosing two flower out of 9 one after the other.
So for say buttercup first flower can be chosen in 3/9 and the next buttercup flower after the fist is removed will be chosen in 2/8.

The same logic applies for the rest of the flowers.
Join the discussion